1. Show that Bohr quantization condition for angular momentum 1 = mur = n? is the same as 2* TT *s= n* i 2. Use uncertain principle ApAr = h /( 2*TT) To show that Minimum radius of Bohrs H-atom with n = 1 4TTE K- rn = n² me2 Hint: express Total energy as a function of r, find the condition for E is minimum. 3. Show that E = (-1/2)* k*e^2/r for H-atom Derive the energy expression as a function of principle quantum number n. E(n) = -((k*e^2/(2r1))(1/n^2) where rn with n = 1(above)
Answer #1

1.

mvr=n\hbar =\frac {nh}{2\pi}

2\pi r =n\frac{ h}{mv}

\lambda =\frac{ h}{mv} (De Broglie relation)

2\pi r =n\lambda

2.

mvr=pr=n\hbar

p =\frac {n\hbar}{r}

E =\frac {n^2\hbar^2}{2mr^2}-\frac {e^2}{4\pi \varepsilon_0 r}

\frac {dE}{dr}= -\frac {n^2\hbar^2 }{mr^3}+\frac {e^2}{4\pi \varepsilon _0 r^2}=0

\frac {n^2\hbar^2 }{mr^3}=\frac {e^2}{4\pi \varepsilon _0 r^2}

r_n=n^2 \frac {4\pi \varepsilon_0 \hbar^2}{me^2}

3.

\frac {mv^2}{r}=\frac {ke^2}{r^2}

mv^2=\frac {ke^2}{r}

E=\frac {1}{2}mv^2 -\frac {ke^2}{r}

E=\frac {1}{2}\frac {ke^2}{r} -\frac {ke^2}{r}

E=\left (-\frac {1}{2} \right )\frac {ke^2}{r}

E=-\frac {1}{2} \frac {ke^2}{n^2 \frac {4\pi \varepsilon_0 \hbar^2}{me^2}}

E=- \frac {m}{ 32\pi \varepsilon_0 \hbar^2n^2}